# Should I Trust You? Detecting Deception in Negotiations using Counterfactual RL

### v2

@@ 6 Results and Analysis
      CTRL-D with human annotated logical forms 0.960 0.300 0.457
@@
Table 3: While LLM baseline Direct Judgment detects deception on actual human lies with a high recall, its precision is very low. LLM baseline using Alignment Judgment and Peskov et al. (2020) LSTM shows problems in detection with poor precision and recall. Our CTRL-D has critically high precision and the best overall tradeoff between precision and recall.
This section compares the approach (Section 3) —a classifier trained by using deceptive values from counterfactual RL and text embeddings from BERT to train a three-layer linear neural network—to an LLM baseline (Section 5). We evaluate these using two Diplomacy datasets; the first is Peskov et al. (2020) with human lie annotations. In Section 6.1, we prove that our approach can outperform the LLM baseline and other baselines in predicting human lies. In Section 6.2, we show that our approach has a high precision in predicting deception, potentially creating friction at just the point when a human player is making a critical decision.
@@ 6.1 Detecting Human Lies
With human lie annotations, we compare CTRL-D prediction to these baselines: 1) LLM baseline using Direct Judgment, 2) LLM baseline using Alignment Judgment and 3) Context LSTM with Power by Peskov et al. (2020) (Table 3). This section answers these questions:
@@
It predicts with high precision $0.950$, compared to both LLM-based methods and to Peskov’s LSTM ($0.147$, $0.095$ and $0.104$). In other words, when CTRL-D predicts a message as a lie, it is 95% likely that the message is an actual human lie. However, our detection is not perfect. Its recall is $0.238$, meaning that it can only catch one out of every four lies. This high precision but low recall is still useful in deployment, since human players at least can trust CTRL-D when it flags lies.
@@
Comparing the recall of CTRL-D with our LLM baseline with Direct Judgment, we find the latter has good recall but low precision. The LLM predicts 45.5% of all messages as lies, which is too sensitive to be useful, when considering the number of misaligned Cicero predicted orders and proposed orders seen in a message. Current LLM-based approaches to providing guidance do not appear capable of recognizing deception, and the resulting high false positive rate makes this strategy not compatible for real-time human games.
@@
To characterize the lies CTRL-D misses detecting (False negatives), we observe that 41 of 61 non-detected lies have no logical forms for negotiations. Further, all of these 41 fit in categories other than Deceptive Moves (Table 2). We examine the remaining 20 and correct errors in logical forms by hand, to see if doing so improves prediction. With logical forms from humans, the recall of CTRL-D (Table 3) improves slightly from $0.238$ to $0.300$, showing that CTRL-D depends on proper logical parsing of natural language text.
@@
We further investigate false positives of CTRL-D and LLM baselines. Since our approach is very precise, it predicts only one true message as a lie (Table 4). On the other hand, the LLM baseline with Direct Judgment [^5] We focus on Direct Judgment and omit the Alignment Judgment baseline variant since both LLM approaches are similar and have similar results. mispredicts $412$ messages as lies (Example in Table 5). We look into these messages and see patterns that LLM baseline is heavily constrained on Cicero’s predicted orders when it considers proposed orders in messages. This could be improved if LLM baseline can recognize the nature of Diplomacy, namely that there are many possibilities for players’ orders, which are not necessarily deceptive.
In sum, CTRL-D captures human lies at best among all methods, including LLM-based methods. Though LLM baseline is better with semantics, it still lacks skills to interpret Diplomacy information in a way that would enable deception detection. With a strong agent, Cicero, predicting human lies using its RL value function makes detection possible. To further validate the quality of our deception detection, we evaluate both CTRL-D and LLM baseline on a larger data set that contains interactions between humans and Cicero.
@@ 6.2 Awareness against Deception
In this section, we evaluate CTRL-D and LLM baselines using the Meta data set. Since this dataset lacks human deception annotations, we first let both models predict whether each message is deceptive, then verify the labeled predictions through human judgment. Human reviewers are provided with historical messages and final orders that sender and recipient submit through games. This information is used to decide whether a sender deceives a recipient by comparing between 1) the sender’s commitment in a proposal and 2) the sender’s final orders. Although this verification is limited to deception that appears within explicit orders of the sender, this could serve as more evidence to verify performance of our approach and the LLM baseline.
CTRL-D predicts with higher precision than LLM baseline, which overpredicts deception.
As shown in Table 6, our findings are consistent with those on the previous dataset. LLM baseline with Direct Judgment predict 41.3% of all samples as deceptive, which is greatly higher than 5% actual lie rate from humans (Peskov et al., 2020). This high rate makes human verification impractical. For LLM baseline with Alignment Judgment, its precision is $0.282$ (only 1 in 4 flagged messages is a true lie). In contrast, CTRL-D, though not matching its previous high performance, is still the best at predicting deception.
@@
We cross-validated our CTRL-D with LLM baseline under Alignment Judgment to evaluate their ability to detect deceptive proposals. While both methods can correctly identify some lies, each may fail under different circumstances. We present several examples here:
@@
7 Related works
Deception in Human Behaviors. Research on deception has highlighted key behavioral and cognitive cues, such as micro-expressions and inconsistencies from mental strain Ekman (2003); Vrij (2008). Multimodal analyses integrating verbal and nonverbal signals have further enhanced detection accuracy DePaulo et al. (2003). Linguistic cues linked to betrayal in the game Diplomacy offer insights closely aligned with our study Niculae et al. (2015). Moreover, computational models using language cues have shown promise in detecting deception in text, though evaluations have been limited to small datasets and specific scenarios Serra-Garcia and Gneezy (2023); Hazra and Majumder (2024). Despite progress, the complex dynamics of deception in human behavior remain underexplored.
ai Deception in Texts. With the rise of ai-generated content, detecting textual deception is crucial. Linguistic and psycholinguistic analysis aids detection, while transformer models improve accuracy Ott et al. (2011); Pérez-Rosas et al. (2015); Zhang et al. (2019). Prior work focuses on detecting ai deception using external and internal methods Park et al. (2024). External techniques like “consistency checks” Fluri et al. (2024) analyze ai behavior for inconsistencies, while internal methods examine embeddings to detect dishonesty Azaria and Mitchell (2023); Burns et al. (2024). Closely related to our work is Fluri et al. (2024), which uses Monte-Carlo Tree Search to check chess moves for logical inconsistencies.
@@ 8 Conclusion and Future Work
Our study confirms that with a well-trained value function, we can estimate deception signs— bait, switch, edge—to predict deception. CTRL-D, our counterfactual RL against deception, has a low recall but a high precision, which can be useful for triggering friction for humans. Comparing to a high recall LLM baseline, CTRL-D is better able to predict human lies and generalize, demonstrating high precision consistently on both evaluation data sets.
While these tasks for deception detection are scoped within Diplomacy, they illustrate the general risks and challenges of ai-deception. Future human-ai interaction tools can build on our methods to reevaluate trust in suspicious negotiations.
@@ Appendix A Lie Categories
  • Feigning Trust/Loyalty. {trust, friend}
  • Withholding Information. {no idea, not sure}
  • False Excuse. {sorry, busy}.

### v3

@@ 6 Results and Analysis
      CTRL-D with annotated logical forms 0.960 0.300 0.457
      CTRL-D with only proposals 0.868 0.413 0.560
@@
Human suspected lies 0.252 0.203 0.225
Table 3: While LLM baseline Direct Judgment detects deception on actual human lies with a high recall, its precision is very low. LLM baseline using Alignment Judgment and Peskov et al. (2020) LSTM shows problems in detection with poor precision and recall. Players from Peskov et al. (2020) struggle to recognize lies, highlighting that our CTRL-D has critically high precision and the best overall between precision and recall.
This section compares the approach (Section 3) —a classifier trained by using deceptive values from counterfactual RL and text embeddings from BERT to train a three-layer linear neural network—to an LLM baseline (Section 5). We evaluate these using two Diplomacy datasets; the first is Peskov et al. (2020) with human lie annotations. In Section 6.1, we prove that our approach can outperform the LLM baseline and other baselines in predicting human lies. Section 6.2 shows that our approach has a high precision in predicting deception, potentially creating friction at just the point when a human player is making a critical decision.
@@ 6.1 Detecting Human Lies
With human lie annotations, Table 3 compares CTRL-D prediction to these baselines: 1) LLM baseline using Direct Judgment, 2) LLM baseline using Alignment Judgment and 3) Context LSTM with power by Peskov et al. (2020) 4) Human suspected lies by Peskov et al. (2020). This section answers these questions:
@@
It predicts with high precision $0.950$, compared to both LLM-based methods and to Peskov’s LSTM. In other words, when CTRL-D predicts a message as a lie, it is 95% likely that the message is an actual human lie. However, our detection is not perfect. Its recall is $0.238$, meaning that it can only catch one out of every four lies. This high-precision but low-recall behavior is still valuable in deployment: since human players alone perform poorly at lie detection (precision = $0.252$), they can rely on CTRL-D’s predictions when it flags a message as a lie.
@@
Comparing the recall of CTRL-D with our LLM baseline with Direct Judgment, the latter has good recall but low precision. The LLM predicts 45.5% of all messages as lies, which is too sensitive to be useful, when considering the number of misaligned Cicero predicted orders and proposed orders seen in a message. Current LLM-based approaches to providing guidance do not appear capable of recognizing deception, and the resulting high false positive rate makes this strategy not compatible for real-time human games.
@@
To characterize the lies CTRL-D misses detecting (False negatives), 41 of 61 non-detected lies have no logical forms for negotiations, while the remaining 20 have errors in logical forms when we parse the messages from English to amr. We examine the remaining 20 and correct errors in logical forms by hand, to see if doing so improves prediction. With logical forms annotated by humans, the recall of CTRL-D (Table 3) improves slightly from $0.238$ to $0.300$, showing that CTRL-D depends on proper logical parsing of natural language text. We further narrow the test dataset to include only messages that explicitly propose moves—allowing us to compute bait, switch, and edge, the three features used to estimate deception. Under this setting, recall increases to $0.413$, resulting in a higher F1-score compared to detecting deception across all messages. While CTRL-D performs best at identifying deceptive proposals involving explicit moves (via reinforcement learning), it also generalizes to detect lies in broader messages through text-based embeddings.
@@
We further investigate the false positives of CTRL-D and LLM baselines. Since our approach is very precise, it predicts only one true message as a lie (Table 4). On the other hand, the LLM baseline with Direct Judgment [^6] We focus on Direct Judgment and omit the Alignment Judgment baseline variant since both LLM approaches are similar and have similar results. mispredicts $412$ messages as lies (Example in Table 5). The LLM baseline is heavily constrained on Cicero’s predicted orders when it considers proposed orders in messages. This could be improved if LLM baseline can recognize that there are many plausible possibilities for players’ orders, which are not necessarily deceptive.
In sum, CTRL-D captures human lies best among all methods, including LLM-based methods. Though LLM baseline is better with semantics, it still lacks skills to interpret Diplomacy information in a way that would enable deception detection. With a strong agent, Cicero, predicting human lies using its RL value function makes detection possible. To further validate the quality of our deception detection, we evaluate both CTRL-D and LLM baseline on a larger data set that contains interactions between humans and Cicero.
@@ 6.2 Awareness against Deception
This section, we evaluate CTRL-D and LLM baselines using the webdiplomacy.net data set. Since this dataset lacks human deception annotations, we first let both models predict whether each message is deceptive, then verify the labeled predictions through human judgment. Human reviewers review with historical messages and final orders that sender and recipient submit through games. This information helps determine whether a sender deceives a recipient by comparing between 1) the sender’s commitment in a proposal and 2) the sender’s final orders. Although this verification is limited to deception that appears within explicit orders of the sender, this could serve as more evidence to verify performance of our approach and the LLM baseline.
CTRL-D has precision than the LLM baseline, which overpredicts deception.
Our findings are consistent with those on the previous dataset, that CTRL-D is the strongest to predict deception (Table 6). LLM baseline with Direct Judgment predict 41.3% of all samples as deceptive, which is greatly higher than 5% actual lie rate from humans (Peskov et al., 2020). This high rate makes human verification impractical. For LLM baseline with Alignment Judgment, its precision is $0.282$ (only 1 in 4 flagged messages is a true lie).
@@
We cross-validate our CTRL-D with LLM baseline under Alignment Judgment to evaluate their ability to detect deceptive proposals. While both methods can correctly identify some lies, each may fail under different circumstances. We present several examples here:
@@
7 Related work
Deception in Human Behaviors. Research on deception highlights key behavioral and cognitive cues, such as micro-expressions and inconsistencies from mental strain Ekman (2003); Vrij (2008). Multimodal analyses integrating verbal and nonverbal signals have further enhanced detection accuracy DePaulo et al. (2003). Linguistic cues linked to betrayal in the game Diplomacy offer insights Niculae et al. (2015). Moreover, computational models using language cues have shown promise in detecting deception in text, though evaluations have been limited to small datasets and specific scenarios Serra-Garcia and Gneezy (2023); Hazra and Majumder (2024). Despite progress, the complex dynamics of deception in human behavior remain underexplored.
ai Deception in Texts. With the rise of ai-generated content, detecting textual deception is crucial. Linguistic and psycholinguistic analysis aids detection, while transformer models improve accuracy Ott et al. (2011); Conroy et al. (2016). Prior work focuses on detecting ai deception using external and internal methods Park et al. (2024). External techniques like “consistency checks” Fluri et al. (2024) analyze ai behavior for inconsistencies, while internal methods examine embeddings to detect dishonesty Azaria and Mitchell (2023); Burns et al. (2024).
@@ 8 Conclusion and Future Work
Our study confirms that with a well-trained value function, we can estimate deception signs—bait, switch, edge—to predict deception. CTRL-D, our counterfactual RL against deceptive proposals, has a good recall and almost perfect precision, which can be helpful for humans that struggle to recognize deception within Diplomacy. Comparing to an LLM baseline that is too sensitive with a higher recall, CTRL-D predicts human lies and generalizes, demonstrating high precision consistently on both evaluation data sets.
While these tasks for deception detection are for Diplomacy, they illustrate the general risks and challenges of ai-deception. Future human-ai interaction tools can build on our methods to reevaluate trust in suspicious negotiations. Applying CTRL-D to real-world deception remains challenging, particularly in open-ended contexts such as scams, where individuals are manipulated into taking actions that compromise their financial security. One possible direction is to extend our method to more controlled environments involving less explicit actions, such as Avalon (Serrino et al., 2019), where deception occurs through proposing team subsets, and eventually to games like Among Us and Werewolves (Jin et al., 2024; Xu et al., 2025), provided there is access to grounded actions, dialogue, and action-value estimations. Looking further ahead, in more complex environments where reinforcement learning has shown strong performance, such as StarCraft and Go (Silver et al., 2016; Vinyals et al., 2019), detecting deception through moves, combat actions, and in-game communication (e.g., language used during battles) presents an exciting direction for future research.
@@ Appendix A Lie Categories
  • Feigning Trust/Loyalty. {trust, friend}
  • Withholding Information. {no idea, not sure}
  • False Excuse. {sorry, busy}.
Appendix B Classifier Discussion
Precision Recall
Rule-based 0.140 0.429
Linear sum with weights 0.250 0.214
CTRL-D (ours) 0.950 0.238
Table A1: Comparison of different deception detection methods.
As there is a concern that a BERT-based classifier may learn a simple heuristic, for example, flagging offers that seem “too good to be true”, rather than truly understanding deception. First, we want to provide evidence that the classifier does not heuristically predict deception. In our preliminary experiments, we first attempted the approach using heuristics directly to deceptive values:
  • Rule-based: $U_{1}>T_{1},U_{2}>T_{2},U_{3}>T_{3}$
  • Linear sum with weights and a threshold: $U_{1}w_{1}+U_{2}w_{2}+U_{3}w_{3}>T$
where results were not good when comparing to the CTRL-D classifier (Table A1).
Precision Recall F1-score
BERT (ours) 0.950 0.238 0.380
RoBERTa (new) 0.767 0.413 0.537
Table A2: Performance comparison of CTRL-D trained on BERT vs RoBERTa.
With a stronger text embedding, e.g. RoBERTa (Liu et al., 2019), we further run additional experiment that replacing BERT with RoBERTa for text embedding and keep the rest of models. The results (Table A2) show RoBERTa can catch more lies. However, we are unsure why precision is slightly lower.
Appendix C Additional Llama experiments
We apply Chain-of-Thought (CoT, Wei et al., 2022) to our LLM baselines. To simplify this, we directly applied CoT into Alignment Judgement, which is one of our LLM baselines that compares the Cicero’s expected action and a dialogue of a certain player to find misalignment that could be a deception. As in Table A6, adding a chain-of-thought by explicitly prompting a text “step-by-step” did not improve from a baseline, Alignment Judgment. We hypothesize that
  1. 1.
    it is still challenging for LLM to recognize patterns from deviations of Cicero’s predicted moves and moves that appear in a conversation in natural language and
  2. 2.
    a possibility of next moves are too wide and Cicero’s predicted moves could not represent a real human intent, relying solely on predicted moves could be impossible.
Another experiment with LLM is providing deceptive values (bait, switch and edge) to our LLM baseline, which results in an LLM triggering lies even more often that its recall is the highest among all baselines; however, the precision of detecting is worsened. Overall, these additional experiments on LLM concretely prove that LLM struggles to detect lies.
@@ Appendix C Additional Llama experiments
Model Precision Recall F1-Score
LLM baseline: Direct Judgment 0.095 0.551 0.161
LLM baseline: Alignment Judgment 0.147 0.065 0.090
LLM baseline: Alignment Judgment and CoT 0.094 0.288 0.141
LLM baseline: bait, switch and edge values 0.081 0.575 0.142
CTRL-D (ours) 0.950 0.238 0.380
Table A6: Additional experiments on prompting Llama with Chain-of-Thought and three deceptive values from our bait, switch and edge estimation.